### v2

Abstract
Accurate rainfall data are crucial for effective climate services, especially in Sub-Saharan Africa, where agriculture heavily depends on rain-fed systems. However, the sparse distribution of rain-gauge networks in the region necessitates reliance on satellite and reanalysis rainfall products (REs) for rainfall estimation. This study evaluated the performance of eight REs — CHIRPS, TAMSAT, CHIRP, ENACTS, ERA5, AgERA5, PERSIANN-CDR, and PERSIANN-CCS-CDR — in Zambia and Ghana using a point-to-pixel validation approach. The analysis encompassed spatial consistency, annual rainfall summaries, seasonal patterns, and rainfall intensity detection across 38 ground stations. Results indicated that no single product performed optimally across all contexts, underscoring the need for application-specific recommendations. All products exhibited a high probability of detection (POD) for dry days in Zambia and northern Ghana (with 70% $<$ POD $<$ 100%, and 60% $<$ POD $<$ 85% respectively), suggesting their potential utility for drought-related studies in these areas. Conversely, all products showed limited skill in detecting heavy and violent rains (with POD close to 0%), rendering them unsuitable for extreme rainfall analysis (such as floods) in the current form. Products integrated with station data (ENACTS, CHIRPS, and TAMSAT) outperformed their counterparts under many contexts, highlighting the importance of calibration with local observations. Bias correction is strongly recommended, as varying levels of biases were evident across different rainfall summaries. A critical area for advancement is extreme rainfall detection. Future research should focus on this aspect.
keywords:
Gauge Observations, Satellite Rainfall Product, Validation, CHIRP, CHIRPS, TAMSAT, ENACTS, ERA5, AgERA5, PERSIANN-CDR, PERSIANN-CCS-CDR
1 Introduction
Accurate historical rainfall observations are crucial for the provision of reliable climate services [2, 6, 17, 52], particularly in regions where agriculture and water resources are highly sensitive to climatic variations [43]. In Sub-Saharan Africa, where about 95 percent of agriculture is rain-fed [1, 49], access to precise rainfall data is essential for effective planning and management of agricultural practices. However, weather station and rain-gauge networks in this region are often sparse, both in spatial and temporal coverage, or in some cases, entirely non-existent [41, 2, 11]. This lack of ground-based observations makes it challenging to generate localized climate information, further complicating efforts to mitigate the impacts of climate variability on agricultural productivity and food security.
REs offer a promising solution by providing rainfall estimates with high spatial and temporal resolutions [21, 47]. These products can potentially enhance climate monitoring and forecasting capabilities in regions with limited ground-based observation networks. However, their accuracy and usability must be rigorously validated against available ground-based observations [17, 6, 2] before they can be effectively integrated into climate services. Validation serves as a key step in assessing how closely these estimates align with actual observations, ensuring their reliability for applications such as agriculture, hydrology, and disaster management [21].
Numerous validation studies have been carried out across the African continent, including those by [19, 13, 17, 37, 2, 21, 47, 59, 45, 30, 31, 6, 61, 24, 63, 36, 46, 26, 44] and [25]. These studies commonly utilize two main approaches for evaluation: pixel-to-pixel and point-to-pixel comparisons [55]. Each approach has distinct advantages and limitations, depending on the specific application and the availability of ground-based observational data.
Pixel-to-pixel validation involves gridding a dense network of rain-gauge stations to a resolution similar to that of the satellite or reanalysis product, allowing for direct comparison across equivalent spatial scales [55, 13, 37]. This method is particularly useful for obtaining a broad, regional assessment of the product’s performance. It is well-suited for applications that require generalized climate information, such as drought monitoring or regional climate modeling. A key advantage of pixel-to-pixel validation is its ability to provide a large-scale comparison that captures spatial rainfall patterns over a wide area. For instance, [13] conducted a pixel-to-pixel validation in Ethiopia using multiple REs, revealing insights into how they perform over various climatic regions. [37] also used pixel-to-pixel validation in East Africa to evaluate the performance of REs, and concluded that while the REs exhibited systematic underestimations, particularly in orographic regions during the October to December rainy season, they could replicate rainfall patterns. One drawback of pixel-to-pixel validation is that it may obscure localized variations, as the averaging effect within grid cells can mask important small-scale rainfall events that are critical for localized agricultural applications.
In contrast, point-to-pixel validation [45, 2, 47, 17, 55] directly compares REs with point-based rain-gauge observations. One advantage of point-to-pixel validation is its ability to provide a more precise assessment at specific locations, helping to identify discrepancies between satellite estimates and on-the-ground observations. However, a key constraint of this approach is the fact that gauges give point measurements while satellites and reanalysis produce spatial averages, which sometimes struggle with local convective storms and orographic rainfall [47]. Comparing rainfall measured at a specific point to satellite estimates averaged over a large area is inherently problematic [47].
In this study, we performed a comprehensive validation of eight REs in Zambia and Ghana using a point-to-pixel approach. This method was chosen primarily because the available station network was sparse, with no more than one station located within a given satellite or reanalysis grid. The uncertainty of interpolation methods increases significantly as the number of stations within a grid decreases [30, 31, 42]. Additionally, we sought to evaluate the suitability of these REs for providing localized climate information. This is due to the growing demand for gridded data with finer spatial and temporal resolutions, driven by the shift in climate change research from global analyses to more regional and localised investigations [30].
Our validation focused on key aspects such as spatial consistency, annual rainfall summaries, seasonal patterns, and rainfall intensity detection. The importance of validating these REs under diverse conditions has been highlighted in previous studies [17]. Through this multi-dimensional approach, we aimed to provide a robust assessment of the utility of these REs for climate services.
The remainder of this paper is structured as follows: Section 2 describes the materials and methods used for the validation, while the results are presented in Section 3, and discussed in Section 4. Finally, the conclusion is given in Section 5. The terms "gauge observations", or "station data", are used interchangeably to refer to rain-gauge station observations or summaries derived from rain-gauge station observations.
2 Materials and Methods
2.1 Study Area
The validation area covers Zambia and Ghana (Figures 1(a) and 1(b) respectively). Zambia lies between 22 to 34 [33] east of Greenwich, and 8 to 18 south of the equator [35]. The prevailing climate is mainly sub-tropical, witnessing around 95 percent of its total precipitation during the wet season extending from October to April [42] when the Inter-tropical Convergence Zone (ITCZ) is located in the region [28]. Zambia is a landlocked, relatively flat country where rainfall is predominantly the result of convection [42]. The rainfall in this region tends to be more uniform and widespread. The variations in rainfall tend to be tied to the specific agroecological region. In the northern areas, the average annual precipitation exceeds 1200 mm/year, while the southern part experiences less than 700 mm/year annually [35]. The central part receives an annual rainfall of approximately 800 - 1000 mm/year, distributed evenly across the crop growing season [33].
Figure 1: Study area
Ghana is located in West Africa, and lies between latitudes 4.6N and 11N, and longitudes -3.3W and 1.8E [50]. The climate of Ghana is tropical and humid [9]. Two rainy seasons occur in the south from April to July and from September to November (with peaks in June and October respectively), whereas the north has only one rainy season from April to September, with one peak in August [3]. Unlike Zambia, the rainfall pattern in Ghana is quite complex with coastal and orographic influences. The seasonality in the rainfall patterns is brought about by the movement of the Inter-Tropical Convergence Zone (ITCZ) [9, 60]. The country can be broadly categorized under three climatic zones, which are the Savannah zone, the Forest zone, and the Coastal zone [8, 50]. Mean annual rainfall ranges between 900 and 2100 mm, where the south-west has relatively high values while the north, and south-eastern coast have relatively low values [5].
The study area was selected based on two key factors:
The first factor was the objective to capture three distinct rainfall patterns across Africa:
  1. 1.
    A unimodal rainfall pattern in the northern part of the continent, with a peak in August, observed in places such as Burkina Faso, Mali, Niger, and northern Ghana.
  2. 2.
    A unimodal rainfall pattern in the southern part of the continent, peaking in February, characteristic of regions like Zambia, Mozambique, and Malawi.
  3. 3.
    A bimodal rainfall pattern in the mid-belt of the continent, found in areas including southern Ghana and parts of Kenya.
By evaluating the performance of REs across these varied rainfall regimes, the study aimed to provide insights into their potential performance in other regions with similar climatic patterns. However, these results are not to be generalized to other locations without validation.
The second factor was the availability of observed rainfall data. We are grateful to the Ghana Meteorological Agency, and the Zambia Meteorological Department for their collaboration and for providing station data for this study to be carried out.
2.2 Data
2.2.1 Station Data
Data from $15$ stations across the country of Zambia, and $23$ stations across Ghana were used in this validation study. The Zambian data was obtained from the Zambia Meteorological Department (ZMD), while the Ghanaian data was obtained from the Ghana Meteorological Agency (GMet). They were daily rain gauge measurements. The data, as were obtained from ZMD and GMet, had longer records going as far back as 1930 for some of the stations. However, 1983 was considered to be the starting period for the stations in order that they coincide with years that most of the REs have available data. Zambian stations with long records (at least 30 years, with the exception of two stations which had 27 and 28 years) were used, and they are scattered unevenly across all the three climatic zones of the country. In the case of Ghana, the 23 stations are synoptic stations [50]. All the data were quality-controlled before use in the study. These quality-control checks included consecutive rain day check, consecutive same rainfall value check, extreme values check, dry months check, and missing values check. This part was mostly done in consultation with the data providers, and took a considerable amount of time as it is a crucial part of the evaluation. Values that did not pass the quality-control checks were replaced with missing data. Table 1 shows the names of the stations, their longitude and latitude information, the period they cover as well as the percentage of complete data within that period.
Table 1: Details of 38 stations in Zambia and Ghana considered for the study
Country Station Latitude Longitude Elevation (m) Period Complete Days (%)
Zambia Chipata -13.64 32.64 1025 1983 - 2022 $98.2$
Zambia Kasama -10.22 31.14 1384 1983 - 2018 $83.7$
Zambia Mansa -11.14 28.87 1257 1983 - 2017 $82.3$
Zambia Mpika -11.90 31.43 1399 1983 - 2020 $83.7$
Zambia Magoye -16.00 27.62 1025 1983 - 2016 $81.4$
Zambia Moorings -16.15 27.32 1085 1983 - 2010 $68.3$
Zambia Choma -16.84 27.07 1275 1983 - 2011 $71.2$
Zambia Livingstone -17.82 25.82 991 1983 - 2019 $89.2$
Zambia Petauke -14.25 31.33 1022 1983 - 2022 $96.4$
Zambia Kaoma -14.80 24.80 1152 1983 - 2022 $83.1$
Zambia Kasempa -13.53 25.85 1134 1983 - 2022 $72.5$
Zambia Mongu -15.25 23.15 1053 1983 - 2021 $95.8$
Zambia Mwinilunga -11.75 24.43 1363 1983 - 2022 $84.0$
Zambia Solwezi -12.18 26.38 1333 1983 - 2022 $99.0$
Zambia Zambezi -13.53 23.11 1078 1983 - 2022 $96.8$
Ghana Abetifi 6.680 -0.747 594.7 1983 - 2022 $92.7$
Ghana Ada 5.778 0.622 5.0 1983 - 2021 91.9
Ghana Akatsi 6.117 0.800 53.6 1983 - 2020 90.7
Ghana Akim Oda 5.929 -0.978 39.4 1983 - 2020 $90.6$
Ghana Akuse 6.095 0.119 17.4 1983 - 2020 94.6
Ghana Axim 4.867 -2.233 37.8 1983 - 2021 93.8
Ghana Babile 10.517 -2.817 304.7 1983 - 2022 $94.8$
Ghana Bole 9.033 -2.483 299.5 1983 - 2022 98.1
Ghana Ho 6.600 0.467 157.6 1983 - 2019 91.9
Ghana Kete-Krachi 7.817 -0.033 122.0 1983 - 2019 $90.0$
Ghana KIAMO-Accra 5.610 -0.168 67.7 1983 - 2021 $95.2$
Ghana Koforidua 6.086 0.271 166.5 1983 - 2022 $97.7$
Ghana Kumasi 6.717 -1.592 286.3 1983 - 2022 $96.6$
Ghana Navrongo 10.878 -1.083 201.3 1983 - 2020 $93.5$
Ghana Saltpond 5.200 -1.067 43.9 1983 - 2019 $91.7$
Ghana Sefwi Bekwai 6.197 -2.321 170.8 1983 - 2021 $96.0$
Ghana Sunyani 7.359 -2.330 308.8 1983 - 2022 $98.3$
Ghana Takoradi 4.894 -1.774 4.6 1983 - 2022 $97.3$
Ghana Tamale 9.554 -0.862 183.3 1983 - 2022 $97.7$
Ghana Tema 5.632 0.002 14.0 1983 - 2021 95.0
Ghana Wa 10.05 -2.50 322.7 1983 - 2020 90.4
Ghana Wenchi 7.75 -2.10 338.9 1983 - 2020 94.2
Ghana Yendi 9.45 -0.03 195.2 1983 - 2021 97.3
[…]
2.2.2 Satellite and Reanalysis Data
Eight REs were used in this study, which include six satellite-based and two reanalysis rainfall products. They are the Climate Hazards Group Infrared Precipitation (CHIRP), CHIRP with Stations (CHIRPS) [23, 22], Tropical Application of Meteorology using Satellite data (TAMSAT) [42, 59, 29], the European Centre for Medium-Range Weather Forecasts (ECMWF) v5 (ERA5) [29, 7], fifth-generation reanalysis of ECMWF (AgERA5 hereafter AGERA5) [10, 12], Precipitation Estimation from Remotely Sensed Information using Artificial Neural Networks-Cloud Classification System-Climate Data Record (PERSIANN-CCS-CDR, hereafter PCCSCDR) [54], Precipitation Estimation from Remotely Sensed Information using Artificial Neural Networks - Climate Data Record (PERSIANN-CDR hereafter PCDR) [4], and Enhancing National Climate Services (ENACTS) [16, 18].
We selected these REs based on the following criteria:
  • Long Historical Records: Each product has a dataset extending over at least 30 years [22, 42], making them reliable for long-term climatic analysis.
  • High Spatial and Temporal Resolution: The REs offer daily or sub-daily data at fine spatial resolutions, which is crucial for capturing localized rainfall patterns.
  • REs which cover the study area, and provide data up to date.
In addition, we also wanted to cover a diverse range of REs that represent different major approaches:
  • Satellite-Based Products: CHIRP, CHIRPS, and TAMSAT have undergone extensive validation in various African regions, and they have been used for a wide range of services on the continent [17].
  • Reanalysis Products: ERA5 [29] is widely recognized as one of the top-performing reanalysis products, with a large body of studies supporting its use across Africa. AGERA5, derived from ERA5, relatively new, is tailored for agricultural and agro-ecological applications [10] but has not been extensively validated, particularly in Zambia and Ghana.
  • Neural Network-Based Products: PCDR and PCCSCDR, both leveraging neural networks, have not been widely validated in the study area, even though PCDR has been recently validated in Côte d’Ivoire [39], parts of west and central Africa [38], Nigeria [51], and Northern Ghana [36].
  • A Met Office-Led Product: ENACTS is a relatively new product that is led by Met offices of different African countries including the Zambia Meteorological Department. A relatively dense network of station data are included in the product, and has a potential added value for Africa [56].
Although other REs, such as the African Rainfall Climatology version 2 (ARC2) [48], the Climate Prediction Center Morphing Method (CMORPH) [34], Integrated Multi-satellitE Retrievals for GPM (IMERG) [53], and the Tropical Rainfall Measuring Mission (TRMM) [57] have been validated in various studies, they failed to meet one or two of the above-mentioned criteria; hence, they were not included in our evaluation. [40] and [55] provide an extensive list of 23 and 44 well-known products respectively with relevant details.
Daily rainfall data from REs were extracted at grid points nearest to the station coordinates (Table 1) for comparison with observations. A summary of the REs’ data used are given in Table 2.
Table 2: Details of satellite and reanalysis rainfall products
Product Inputs Spatial Coverage Period Spatial Resolution Temporal Resolution
CHIRPS Satellite + gauge merge Global 1981 - present 0.05 Daily
CHIRP Satellite Global 1981 - present 0.05 Daily
TAMSAT Satellite + gauge calibration Africa 1983 - Present 0.0375 Daily
ERA5 Reanalysis Global 1940 - present 0.25 Hourly
AgERA5 ERA5 Global 1979 - present 0.1 Daily
ENACTS Satellite + gauge merge Selected countries (including Zambia) 1981 (for Zambia) - present 0.0375 Daily
PCDR Satellite Global 1983 - present 0.25 Daily
PCCSCDR Satellite Global 1983 - present 0.04 Daily
[…]
For REs that incorporate station data, such as CHIRPS, some studies exclude stations used in the product’s development during the validation process to ensure an unbiased comparison [17]. While we acknowledge this approach as methodologically sound, in this study, we chose not to exclude any stations. Our rationale is twofold: first, we prioritized the perspective of end users, who are unlikely to know which stations are integrated into the products and simply seek guidance on which product is most suitable for their location and/or application of interest. Second, by including all stations in the validation, we gain insight into the product’s performance both at locations used in its development and at those that were not. This approach allows us to assess the product’s generalizability across unseen or arbitrary locations, providing a more comprehensive evaluation. While this was not the primary aim of our analysis, it is highlighted at specific points where it is particularly relevant. Third, many researchers compute summary statistics (e.g., mean bias, root mean squared error) across all available stations for each product to facilitate easy comparison and ranking. However, this approach can introduce bias if stations used in the product’s development are included in the validation process, potentially inflating performance metrics [17, 61]. Our study took a different approach; instead of focusing on the overall performance across all stations, we were interested in evaluating how each product performs at individual locations. While this method can be more labor-intensive, especially when working with a dense network of stations, it provides a more granular understanding of product accuracy and suitability for localized climate applications. This site-specific focus ensures that we assess each product’s performance without the bias introduced by averaging across multiple stations.
The ERA5 hourly values were accumulated to daily values. In the course of this study, PCCSCDR was no longer available for download and could not be downloaded for Ghana, and hence it was not considered in the validation for Ghana. Part of the reasons for that may be due to spatial consistency check in Section 2.3 and Section 3.1.
2.3 Checking Spatial Consistency of REs
The spatial consistency of REs was examined, focusing on annual total rainfall, rain day frequency, and mean rain per rainy day. The interest was to ensure uniform spatial variation in these summaries, without station data comparison. Maps of the long-term averages of these summaries were used to study how well the climatology of Zambia and Ghana is captured by the REs, allowing for detection of any unnatural spatial variations.
2.4 Comparison of Annual Summaries
Annual summaries (total rainfall, number of rainy days, and mean rain per rainy day) were used to study this aspect of the REs. In the case of Zambia, each year was adjusted to commence in August and end in July in order to capture the whole rain and dry seasons in one cycle.
These summaries were then compared with the observed rainfall summaries. Mean error (ME), percentage bias (PBIAS), the ratio of standard deviations (RSD), and linear correlation coefficient ($r$) were used as continuous evaluation metrics. The ME (1) was used to measure the average difference between REs and gauge observations, the PBIAS (2) was used to assess the systematic bias of REs [41], the r (3) to assess the degree of linear relationship between the REs and gauge observations [62], and the RSD (4) to compare the similarity between the variabilities of the REs and the observed rain rainfall data.
$\text{ME}=\frac{1}{N}\sum_{i=1}^{N}(S_{i}-O_{i})$ (1)
$\text{PBIAS}=\frac{\sum_{i=1}^{N}(S_{i}-O_{i})}{\sum_{i=1}^{N}O_{i}}\times 100$ (2)
$r=\frac{\sum_{i=1}^{N}({S}_{i}-\bar{S})({O}_{i}-\bar{O})}{\sqrt{\sum_{i=1}^{N}({S}_{i}-\bar{S})^{2}\sum_{i=1}^{N}({O}_{i}-\bar{O})^{2}}}$ (3)
$\text{RSD}=\frac{{\sigma}_{s}}{{\sigma}_{o}}$ (4)
where $S_{i}$ is the RE series, $O_{i}$ is the gauge observation series, $N$ is the number of data pairs, $\sigma_{s}$ and $\sigma_{o}$ are the standard deviations of the RE and gauge observation series respectively, $\bar{S}$ and $\bar{O}$ are the means of the RE and gauge observation series respectively.
2.5 Comparison of Seasonal Behaviour
To analyze rainfall frequency and intensity, we employed Markov chain models [60], focusing on the Zero-Order Markov chain to estimate daily rainfall occurrence. The rainfall occurrence was modeled as a proportion for each day of the year using logistic regression, with Fourier series applied to capture the periodic nature of seasonal variation. This approach allowed for a detailed comparison between REs and gauge observations, facilitating an evaluation of rainfall frequency patterns across the annual cycle.
For each of the 15 stations in Zambia and the 23 stations in Ghana, separate zero-order Markov chain models incorporating three harmonics were fitted to both the gauge observations and the REs. The model formula is represented by (5) below:
$y(t)=\beta_{0}+\sum_{i=1}^{k}\Big{[}A_{i}\cos\left(\frac{2\pi it}{p}\right)+B_{i}\sin\left(\frac{2\pi it}{p}\right)\Big{]}+\epsilon,$ (5)
where $k$ is the number of harmonics, $p$ represents the period, $t$ is the time (day of year), $\beta_{0}$, $A_{i}$, and $B_{i}$ are the model parameters, $\epsilon$ is the error term, and
$y(t)=\begin{cases}0&\text{if }x(t)<Tr\\ 1&\text{Otherwise}\end{cases}$ (6)
where $x(t)$ is the rainfall value on day $t$ in mm, and $Tr$ is the rainy day threshold in mm.
Importantly, only days where both gauge and product data were available were used in the models to ensure direct comparability.
Additionally, we explored the sensitivity of the results to varying rainy day thresholds. Separate models were developed for $Tr=$ 0.85 mm, 2 mm, 3 mm, 4 mm, and 5 mm for each product (while keeping the $Tr$ of the gauge at 0.85 mm) to examine whether adjusting rain day thresholds could mitigate biases. This approach enabled an assessment of how these biases and threshold adjustments might vary by location and time of year.
2.6 Rainfall Intensity
In order to understand the ability of the REs to detect different types of rainfall, they were assessed over a set of rainfall intensity categories. The categories are defined in (7).
The following terms were also used:
  • True Positive (Hit): the number of observed rainfall events correctly detected by the RE.
  • False Positive (False Alarm): the number of instances where the RE detected a rainfall event that was not observed.
  • False Negative (Miss): the number of events where a rainfall event was observed but was not detected by the RE.
  • True Negative: the number of instances where the RE correctly detects a dry day as observed.
$f(x)=\begin{cases}\text{Dry}&\text{if }x<0.85\\ \text{Light Rain}&\text{if }0.85\leq x<5\\ \text{Moderate Rain}&\text{if }5\leq x<20\\ \text{Heavy Rain}&\text{if }20\leq x<40\\ \text{Violent Rain}&\text{if }x\geq 40\end{cases}$ (7)
where $x$ is the daily rainfall value in mm. The intensity category definitions are similar to that of [64] except the dry day threshold is slightly lower based on reason by [58].
With rainfall intensity classification, we used probability of detection, which is a standard measure for binary classification. It is the likelihood that the REs will correctly detect the same rainfall events on the same days as recorded in the gauge observations. The probability of detection (POD) is given by (8).
$\text{POD}=\frac{\text{Hit}}{\text{Hit}+\text{Miss}}$ (8)
3 Results
The results from the study are presented and discussed under this section.
3.1 Spatial Consistency of REs
For the case of Zambia, it was found that, with the exception of PCCSCDR, none of the REs exhibited any form of spatial irregularities for mean annual total rainfall, mean annual number of rainy days, and mean rain per rainy day. This is an indication that these REs most likely effectively capture Zambia’s rainfall patterns. For instance, Figure 2, showing the spatial map of CHIRPS for mean annual total rainfall, reveals pronounced annual rainfall in the northern region (consistent with the climatology of Zambia), and lacking any noticeable unnatural variability. Consequently, rainfall estimates from nearby weather stations are likely to align closely. Applications relying on these summaries are likely to align well with gauge observations.
Figure 2: Mean total rainfall (mm/year) of CHIRPS for 1983-2022 in Zambia
On the contrary, the mean annual total rainfall from the PCCSCDR product exhibits spatial irregularities characterized by localized pixelated variations, with sharp contrasts in colouration at the boundaries of the squares compared to the more uniform colouration within them, as shown in Figure 3. These contrasts suggest that extracting information from the dataset may yield significantly different values depending on whether the data is sourced from the boundaries or within the squares, despite their close proximity. This spatial inconsistency implies that variations in the data may arise not from actual geographical differences but rather from the pixelation effect. Consequently, such inconsistencies could lead to misleading interpretations in applications that rely on accurate spatial rainfall distribution.
Figure 3: Mean total rainfall (mm/year) of PCCSCDR for 1983-2022 in Ghana
In response to the spatial inconsistency observed from PCCSCDR, we contacted the team of researchers at the Center for Hydrometeorology and Remote Sensing (CHRS) at the University of California, Irvine (UCI). They acknowledged recent observations of potential issues with the product and were investigating the underlying causes of these inconsistencies. Consequently, PCCSCDR was excluded from further consideration in this study. There are no noticeable irregularities from the other REs in Zambia, and can be seen from Figure 4.
Figure 4: Mean total rainfall (mm/year) of the REs for 1983-2022 in Zambia
Figure 5: Mean total rainfall (mm/year) of the REs for 1983-2022 in Ghana
In the case of Ghana, none of the REs displayed spatial inconsistencies and were found to align closely with the country’s climatology. Figure 5c illustrates the mean annual total rainfall for PCDR, which exhibits a pixelated appearance resulting from its coarse resolution of 0.25 by 0.25 (approximately 25 km by 25 km). This pixelation is similar to that observed in ERA5 (and AGERA5 derived from ERA5), as shown in Figure 5a, while CHIRPS and CHIRP (Figures 5b and 5d, respectively) feature a finer resolution of 0.05 by 0.05 (approximately 5 km by 5 km). Despite the pixelation, all REs effectively capture the climatological patterns in Ghana, with the northern regions and the east coast receiving less rainfall, and the central belt and west coast (particularly the Axim area) experiencing significantly higher precipitation. Overall, ENACTS, CHIRPS, CHIRP, and TAMSAT demonstrated a superior spatial consistency with the climatology of Zambia and Ghana.
3.2 Comparison of Annual Summaries
The performance of the REs on annual summaries (number of rainy days, total rainfall, and mean rain per rainy day) are presented below:
3.2.1 Number of rainy days/rain day frequency
The line plots in Figures 6 and 7show the MEs on the number of rainy days of each product at the various stations in Zambia and Ghana respectively. The height of the bars at the different stations represents the mean number of rainy days at those stations. Each bar is divided into four equal segments (separated by horizontal white lines), with the first segment at the bottom and the fourth segment at the top. These divisions facilitate the estimation of PBIAS. For example, TAMSAT has an ME of about 40 mm at Chipata while having a PBIAS of about 50% (see Figure 6).
Figure 6: ME of number of rainy days in Zambia for each RE on the quartiles of the number of rainy days at each station
Figure 7: ME of number of rainy days in Ghana for each RE on the quartiles of the number of rainy days at each station
On average, all the REs overestimated the number of rainy days in Zambia and Ghana (Figures 6 and 7).
Among these products, ENACTS showed the lowest average bias in Zambia, followed by CHIRPS and TAMSAT. Apart from ENACTS, CHIRPS, and TAMSAT, the remaining products exhibited an average bias exceeding 50% in Zambia, indicating a considerable overestimation of rainy days.
The overestimation was higher in Ghana. With the exception of CHIRPS and TAMSAT, the remaining REs exhibited an average bias exceeding 100% (Figure 7). For CHIRPS, TAMSAT, and, to a lesser extent, PCDR, the overestimation tended to be more pronounced in the northern stations (Tamale, Wa, Yendi, Babile, Bole, and Navrongo) than at the southern stations. On the contrary, the overestimation by the other REs was more pronounced, on average, at the southern stations than the northern stations.
Figures 8 and 9 illustrate the performs of the REs in detecting rainy days across various stations in Zambia and Ghana respectively.
Figure 8: Stacked bar chart illustrating the performance of different REs in detecting rainy days across various stations in Zambia, as represented in the graph facets.
Figure 9: A stacked bar chart showing how well the REs detect rainy days at the different stations in Ghana (shown in different facets of the graph)
From Figures 8 and 9, it was observed that all the products had considerable proportions of False Positives. This occurrence was lowest in ENACTS, followed by CHIRPS and TAMSAT. Proportion of False Negatives were lowest in ERA5, AGERA5, and PCDR and was relatively higher in CHIRPS, ENACTS, and CHIRP.
3.2.2 Annual rainfall totals
As compared to rainfall frequency, the biases in total rainfall were generally low for all the REs both in Zambia and Ghana ($-30$% $\leq$ PBIAS $\leq$ 30%) as can be seen in Figures 10 and 11 respectively. CHIRPS, TAMSAT and CHIRP had the least biases ($-20$% $<$ PBIAS $\leq$ 10%) in Zambia. ENACTS, CHIRPS, CHIRP, and TAMSAT underestimated total annual rainfall at most of the stations in Zambia while ERA5 and AGERA5, overestimated total rainfall at most of the stations in Zambia (Figure 10). PCDR overestimated at some of the stations (0% $<$ PBIAS $<$ 20%) while underestimating at some (with $-20$% $<$ PBIAS $<$ 0%). The product with the highest underestimation was ENACTS ($-30$% $\leq$ PBIAS $<$ 0%) across all stations while AGERA5 was the product with the highest overestimation (0% $<$ PBIAS $\leq$ 30%) at most of the stations.
Figure 10: Line plots showing PBIAS, r, and RSD on annual total rainfall for all the products (in different colours and line types) across all the stations (on the x-axis) in Zambia. The y-axis displays the values of the corresponding metric
Figure 11: Line plots showing PBIAS, r, and RSD on annual total rainfall for all the products (in different colours and line types) across all the stations (on the x-axis) in Ghana. The y-axis displays the values of the corresponding metric
The total rainfall of the REs tended to correlate well with observed data (0.5 $<$ r $<$ 1) across most of the stations in Zambia (Figure 10). ENACTS seemed to have the highest correlations across the stations, followed by CHIRPS. CHIRP appeared to have the lowest correlations, generally less than 0.6. While CHIRPS and TAMSAT appeared to correlate well with gauge observations in Ghana (Figure 11), with 0.5 $<$ r $<$ 1, the correlation from the other products were not as good (r $<$ 0.6 at most of the stations). ERA5 had the lowest correlations, followed by AGERA5 and CHIRP.
The REs tended to have less variability than the observed rainfall total (RSD $<$ 1) across most of the stations in Zambia and Ghana (Figures 10 and 11 respectively).
3.2.3 Mean rain per rainy day
Figure 12: Line plots showing PBIAS, r, and RSD on mean rain per rainy day for all the products (in different colours and line types) across all the stations (on the x-axis) in Zambia. The y-axis displays the values of the corresponding metric
Figure 13: Line plots showing PBIAS, r, and RSD on mean rain per rainy day for all the products (in different colours and line types) across all the stations (on the x-axis) in Ghana. The y-axis displays the values of the corresponding metric
The REs underestimated mean rain per rainy day at varying degrees, with $-60$% $<$ PBIAS $<$ 0% at most of the stations in Zambia (Figure 12), and $-75$% $<$ PBIAS $<$ 0 at most of the stations in Ghana (Figure 13). Correlations were generally low for most of the products across all the stations. Aside from ENACTS (with 0.5 $<$ r $<$ 1 at most of the locations), most of the REs had correlations between $-0.3$ and 0.5 at most of the stations in Zambia and Ghana (respectively Figures 12 and 13). The REs also tended to have lesser variability (RSD $<$ 0.8) across most of the stations as compared to gauge observations in Zambia and Ghana.
3.3 Comparison of Seasonal Behaviour
Figures 14-20 show the Zero-order Markov chain models with three harmonics fitted separately to the rainfall occurrence at each of the stations in Zambia and Ghana for the gauge observations as well as the REs. At each station, only days which the gauge observations and RE values were not missing were included to ensure comparison of the same days. The solid black curve shows the fitted rain day frequency for the gauge observations with 0.85mm rain day threshold. The coloured dashed lines show the fitted rain day frequency for the RE values at various rain day thresholds (0.85mm, 2mm, 3mm, 4mm, and 5mm). The y-axis on each plot represents the proportion of rain day frequency, while the x-axis shows the date in the year, spanning from August 1 to July 31 for Zambia, and January 1 to December 31 for Ghana.
Figure 14: Rain day frequency models of gauge observations vs ERA5 in Zambia
Figure 15: Rain day frequency models of gauge observations vs TAMSAT in Zambia
Figure 16: Rain day frequency models of gauge observations vs CHIRPS in Zambia
Figure 17: Rain day frequency models of gauge observations vs ENACTS in Zambia
Figure 18: Rain day frequency models of gauge observations vs PCDR in Ghana
Figure 19: Rain day frequency models of gauge observations vs TAMSAT in Ghana
Figure 20: Rain day frequency models of gauge observations vs CHIRPS in Ghana
The seasonality at the Zambian stations were well captured by all the REs, with the highest rain day frequency in the November to May season (see Figures 14-17 for ERA5, TAMSAT, CHIRPS, and ENACTS respectively consistent with the REs not shown here).
At the 0.85 mm rain day threshold, the REs overestimated the rain day frequency in varying degrees. In general, ENACTS seemed to have the lowest overestimation, followed by CHIRPS and TAMSAT. ERA5 (and AGERA5, PCDR, CHIRP, not shown here) had a higher degree of overestimation.
Most of the REs tended to match the gauge observations better at a higher rain day frequency, while still capturing well the seasonality.
There was an indication that the choice of threshold may depend on the the time of year. For example, the choice of 5 mm for ERA5 rain day threshold seemed to be optimal for most parts of the year at Mpika, but was not optimal between January and March (see Figure 14).
There was also an indication that the choice of threshold may depend on the location. For instance, the choice of 5 mm rain day threshold seemed optimal for ERA5 at Solwezi, but not at Zambezi (see Figure 14).
Last but not least, the choice of threshold also appeared to be dependent on the product. For example, while the choice of 5 mm rain day threshold appeared to be optimal for TAMSAT at Magoye (Figure 15), it was not optimal for ENACTS at the same station (Figure 17).
Similar to the case of Zambia, all the REs captured the seasonality well at the southern stations of Ghana, capturing the bi-modal rainfall patterns (April-July, and September-November, with peaks around June and October respectively).
However, at 0.85mm threshold, CHIRPS overestimated rain day frequency between February and June while it underestimated rain day frequency between June and October at Abetifi, Akim Oda, Akuse, Axim, Ho, KIAMO Accra, Koforidua, Kumasi, Saltpond, Sefwi Bekwai, Sunyani, and Takoradi (Figure 20). TAMSAT also underestimated rainday frequency between June and October at Akim Oda, Axim, Koforidua, Sefwi Bekwai, and Takoradi (Figure 19). While increasing thresholds seemed to fix the overestimation for the other REs across all stations, the underestimation by CHIRPS and TAMSAT did not seem to change under different rain day thresholds. There is potentially an issue with these REs at these locations.
Aside from CHIRPS and PCDR, the REs also captured seasonality of the northern stations of Ghana well, capturing the unimodal patterns. CHIRPS and PCDR (Figures 20 and 18 respectively) exhibited bi-modal patterns at these northern stations, which have unimodal rainfall patterns.
Similar to Zambia, most of the REs tended to match the gauge observations better at a higher rain day frequency, while still capturing well the seasonality. These thresholds seemed to depend on the RE, location and time of year (similar to the case of Zambia).
3.4 Rainfall Intensity
Figure 21 and 22 show the different REs and their probability of detection of different rainfall intensity categories at the different stations in Zambia and Ghana respectively.
Figure 21: Probability of detection for different rainfall intensity categories in Zambia
Figure 22: Probability of detection for different rainfall intensity categories in Ghana
All the REs showed high skill of detecting dry days (with 70% $<$ POD $<$ 100%) across all stations in Zambia (Figure 21). ENACTS, CHIRPS and TAMSAT had about 10% chance higher in detecting dry days than the other REs. Apart from CHIRP, all the REs had POD $>$ 50% in detecting moderate rain at almost all the stations, with ENACTS having 50% $<$ POD $<$ 80%, outperforming all the REs, followed by TAMSAT (with 50% $<$ POD $<$ 70%). All the REs had a probability of less than 50% in detecting light rain. They all performed poorly in detecting heavy and violent rains (with POD close to zero across all stations).
In the case of Ghana, not all the REs demonstrated a high probability of accurately detecting dry days, as shown in Figure 22. Aside from TAMSAT and CHIRPS, which exhibited a POD of at least 70% across all stations, the other REs had POD often below 70% in southern Ghana, and 60% $<$ POD $<$ 85% in northern Ghana. The two reanalysis products — ERA5 and AGERA5 — stood out slightly in the detection of light rain, even though they both had a POD below 50% at most of the stations. TAMSAT outperformed all the REs in the detection of moderate rain (with 50% $<$ POD $<$ 80%) at almost all the stations. The worst performance of most of the REs in detecting dry days was observed in stations located in the central and southern regions, particularly along the southwestern coast of the country.
Similar to Zambia, the REs performed poorly in the detection of heavy and violent rains (with POD close to zero across all stations).
Figure 23: Graph of Kasama station rainfall intensity categories
Table 3: Percentages of observed rainfall intensity categories at the various stations in Zambia
Station Light Rain (%) Moderate Rain (%) Heavy Rain (%) Violent Rain (%)
Chipata 36 39 16 9
Choma 40 43 12 5
Kaoma 42 41 13 5
Kasama 37 42 16 5
Kasempa 38 42 15 5
Livingstone 43 40 12 5
Magoye 41 40 14 5
Mansa 37 43 15 6
Mongu 39 42 14 5
Moorings 34 46 14 6
Mpika 41 40 15 5
Mwinilunga 41 41 13 4
Petauke 39 39 16 6
Solwezi 36 43 16 5
Zambezi 40 42 14 5
[…]
Figure 23 represents the distribution of hits and misses in the different rainfall intensity categories at Kasama, Zambia. True Hit represents the proportion of rainfall events in the rainfall intensity category that were correctly detected by the RE; True Miss represents the proportion of rainfall events in the rainfall intensity category that the RE detected as a dry day. Lower represents the proportion of rainfall events in the rainfall intensity category that the RE missed by estimating a lower intensity category. Higher represents the proportion of rainfall events in the rainfall intensity category that the RE missed by estimating a higher intensity category. The distribution of hits and misses in Kasama was not very different from those at other stations in Zambia and Ghana (not shown). True Hit was relatively high for moderate rain, which relates to the observation in Figure 21. As expected, True Miss was very low across all intensity categories. True Miss was the lowest under violent rain followed by heavy rain. The figure further shows that rainfall was correctly detected by most of the REs on days of violent and heavy rain, but the intensity was usually rather underestimated.
Table 3 also gives percentages of different rainfall intensity category observed at the various stations in Zambia. This table shows that heavy and violent rains were usually less than 20% and 10% respectively of all observed rainfall rainfall events across all stations.
4 Discussion
The results have shown that no single product is universally optimal across all contexts highlighting the fact that product recommendations must be specific to the intended application, addressing questions such as:
  1. 1.
    Can the product reliably monitor extreme rainfall events in region X?
  2. 2.
    Is it suitable for drought analysis in region Y?
  3. 3.
    Does the product adequately capture seasonality in location Z?
This targeted validation enables more informed and context-specific decision-making.
The topography in the southern part of Ghana (including locations like Axim and Takoradi) is complex due to the vegetation and terrain, and its proximity to the coast and surrounding forest regions. These regions are subject to localized weather patterns [3], which can be harder for satellite-based or reanalysis products to accurately capture. Coastal influences, such as sea breezes, can induce small-scale convective systems that could be missed by the REs. On the contrary, Zambia is predominantly flat and landlocked, with rainfall largely driven by convective systems, and there is typically a strong correlation between the cold cloud tops of these systems and rainfall [42]. Our results confirm the fact performance of the REs under annual summaries, seasonality studies, and rainfall intensity detection was more consistent in Zambia as compared to Ghana.
Previous works have shown that REs are usually good at detecting dry days at many places [46, 64]. From our results, all the REs were seen to detect dry days well in Zambia and northern Ghana while CHIRPS and TAMSAT still maintained consistent high performance in south Ghana as well, consistent with previous existing literature. One explanation could be as explained in the previous paragraph. Another reason could be due the fact that TAMSAT and CHIRPS have been designed for drought monitoring [22, 42].
It is known that the REs are extensively validated on their ability to capture daily, sub-seasonal, seasonal, or annual rainfall amounts during development [22, 42, 4]. Our results shown on Figures 10 and 11 show low biases and good correlations on total rainfall, consistent with [2], and likely explained by the literature above.
All the REs capture the seasonality well in both countries at the 0.85mm (or higher, or both) rain day threshold. The good performance of the REs at the seasonal scale is in agreement with results by other researchers [2].
ENACTS includes a large number of meteorological station data than other similar blended REs [56, 18]. CHIRPS also includes some station data, while TAMSAT is calibrated with station data. Our results showed that ENACTS stood out as one of the best products across multiple contexts in Zambia (followed by CHIRPS and TAMSAT). This is likely due to the merging/calibration of these products with station data. Local calibration could be a potential approach for reducing the satellite precipitation errors, as was also noted by [14] and [15].
REs have been known with their tendency to overestimate the number of rainy days, as confirmed by previous works within the Africa region [44]. Our results as shown in Figures 6 and 7 is consistent with existing findings. This tendency to overestimate rainy days is expected, as the REs represent area-based precipitation estimates, while gauge observations are point measurements. Due to the overestimation, they had low instances of False Negatives. This is confirmed by the results in Figures 8 and 9 for Zambia and Ghana respectively. They also had considerable proportions of False Positives due to the overestimation (see Figures 8 and 9). Some of these False Positives may have stemmed from rainfall occurring elsewhere within the grid area but outside the immediate vicinity of the gauge. Further investigation, using additional station data from within these grids, could clarify these discrepancies. It is also possible that rainfall detected by REs occurs several kilometers away or even on a different day, as noted by [32]. Rain-bearing clouds can travel hundreds of kilometers before releasing rain, sometimes accumulating over several days before delivering intense rainfall in a single event [27].
All the REs showed poor performance in detecting heavy and violent rainfalls (see Figures 21 and 22), which is consistent with existing literature [62, 46, 64, 2]. This indicates that the products in their current state are likely unsuitable for extreme rainfall detection or studies at these locations. Developing products tailored to extreme rainfall detection, which could then complement existing products, is imperative. [20] noted that these extreme events have not been extensively explored. As illustrated in Figure 23, most heavy and violent rain days were correctly classified as rainy by the products, albeit with underestimated rainfall amounts. This suggests that focused enhancements in design could significantly improve detection accuracy and rainfall estimation for these events.
5 Conclusion
This study evaluated the performance of eight satellite and reanalysis rainfall products — CHIRPS, TAMSAT, CHIRP, ENACTS, ERA5, AgERA5, PERSIANN-CDR, and PERSIANN-CCS-CDR — in Zambia and Ghana using a point-to-pixel validation approach. The analysis encompassed spatial consistency, annual rainfall summaries, seasonal patterns, and rainfall intensity detection across 38 ground stations. The results have provided useful insights into the performance of the products and the implication on their usage as follows:
All products exhibited high probability of detection (POD) for dry days in Zambia and nothern Ghana (with 70% $<$ POD $<$ 100%, and 60% $<$ POD $<$ 85% respectively) while CHIRPS and TAMSAT maintained a consistent high performance in southern Ghana. These products may potentially be useful for drought detection at those locations.
There were biases in the products under multiple summaries (including total rainfall, rain day frequency). Bias-correction is recommended, especially for reducing the biases on the number of rainy days.
The products merged/calibrated with station data (ENACTS, CHIRPS, and TAMSAT) appeared to perform better than the other REs under many of the contexts. This is perhaps one potential route for improving the performance of REs.
Lastly, all REs performed poorly in the detection of heavy and violent rainfall (with POD close to 0%). At their current state, the products may not be recommended for extreme rainfall events detection at these locations, such as floods. Future research should look more into this.
[…]
Supplementary information
Not applicable
[…]
Acknowledgments
The authors acknowledge the Ghana Meteorological Department, the Zambia Meteorological Department for their collaboration and making station data available for this study.
Declarations
  • Funding This publication was made possible by a grant from Carnegie Corporation of New York (provided through the African Institute for Mathematical Sciences). The statements made and views expressed are solely the responsibility of the author(s).
  • Competing interests The authors have no relevant financial or non-financial interests to disclose.
  • Ethics approval Not applicable
  • Consent to participate Not applicable
  • Consent for publication Not applicable
  • Data Availability The RE data used are all publicly available. The station data for Ghana and Zambia can be obtained from GMet and ZMD respectively.
  • Code availability Code is available upon request from first author.
  • Authors’ contributions Conceptualization: J.B., D.S.; Methodology: J.B., D.S.; Formal analysis: J.B., D.S.; Data curation: J.B., D.S., F.F.T.; Writing — original draft: J.B.; Writing — review & editing: J.B., D.S., D.N., F.F.T.; Visualization: J.B., D.S.; Supervision: D.S., D.N., F.F.T.